OpenStudy (anonymous):

Convergent or divergent? \[\sum_{n = 1}^{\infty} \frac{n!}{e^{n!}}\]

OpenStudy (anonymous):

this series converges by Ratio test..

OpenStudy (anonymous):

hmmm did you studied taylor series?

OpenStudy (anonymous):

yes.. why?

OpenStudy (anonymous):

it approaches to zero

OpenStudy (anonymous):

\[e^{n!}>>n!\]

OpenStudy (anonymous):

that reasoning may not work sometimes. consider:\[\sum_{n = 1}^{\infty} \frac{1}{n}\] 1>>n as n -> infinity but it would still diverge.

OpenStudy (anonymous):

Sorry thats 1 << n

OpenStudy (anonymous):

The ratio test should be used I think \[\left|\frac{u_{n+1}}{u_n}\right| = \left|\frac{(n+1)!e^{n!}}{n!e^{(n+1)!}}\right| = \left|\frac{(n+1)}{e^{n!n}}\right|\] since \[e^{n!}/e^{(n+1)!} = e^{n!}/(e^{n!})^{(n+1)} = 1/(e^{n!})^n\] so \[\lim_{n\rightarrow \infty}\left|\frac{u_{n+1}}{u_n}\right| = 0 < 1\] so the sum converges.

OpenStudy (anonymous):

there is mistake \[\left| \frac{n+1}{e^{n+1}} \right|\]